OpenStudy (anonymous):

What is 2^(4/3) equal to

OpenStudy (studygurl14):

Are you looking for the actual calulation of it? or the conversion into a radical?

OpenStudy (anonymous):

cube root of 8 square root of 8 cube root of 16 square root of 16

OpenStudy (anonymous):

Like those^

OpenStudy (studygurl14):

radical...ok hold on

OpenStudy (studygurl14):

\[2^{4/3} = \sqrt[3]{2^4}=\sqrt[3]{16}\]

OpenStudy (studygurl14):

so C

OpenStudy (studygurl14):

http://www.mathsisfun.com/algebra/exponent-laws.html

OpenStudy (anonymous):

I thought so, thanks.

OpenStudy (studygurl14):

yw